An object is projected directly up so that its height in metres at time t seconds can be modelled by

h(t) = -0.5t2 + 9t + 9.1

a. From what height was the object initially projected?

b. What was the initial velocity?

c. Find the velocity and height when t = 9s.

d. When does the object return to its initial height?


1
Expert's answer
2021-02-28T16:08:12-0500

The height of the object in meters at time "t" seconds is given by "h(t)=-0.5t^2+9t+9.1"

(a) At the initial height "t=0."

"\therefore h(0)=9.1"

Therefore the object initially projected from the height "9.1" meters.

(b) Velocity "=" Rate of change of height "=" "\frac{dh}{dt}" .

"\therefore \frac{dh}{dt}=[(-0.5)×2t]+9=(-t+9)"

For the initial Velocity "t=0."

"\therefore [\frac{dh}{dt}] _{t=0}=9"

Therefore the initial Velocity of the object is "9" "m\/sec."

(c) At "t=9s" , "h(9)=[(-0.5)\u00d79^2]+(9\u00d79)+9.1"

"=(-40.5)+81+9.1"

"=49.6"

and "[\frac{dh}{dt}]_{t=9}= (-9+9)=0"

Therefore at "t=9s" , the height of the object is "49.6" meters and Velocity is "0."

(d) The initial height of the object is "9.1" meters.

According to the problem, "9.1=(-0.5)t^2+9t+9.1"

"\implies 0.5t^2=9t"

"\implies t^2-18t=0"

"\implies t=0" and "t=18"

But "t=0" is the initial position of the object.

So the object return to its initial height at "t=18" "sec."